Shift report 20000715 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sat Jul 15 2000 - 07:59:26 EDT

  • Next message: http account: "Shift report 20000715 08:00-16:00"

    ------------------------------------------------------------
    Date: Sat Jul 15 07:59:26 2000, Shift: 08:00-16:00
    Supervisor: Michael Murray, with: Sorin Zgura
    Log book#3 pages 9 - 16
    
    Shift Summary:
    ------------------------------------------------------------
    The RHIC group worked on injection most of the night.
    Ussually there was only 1 beam or none.
    
    Runs 1784,1785, 1786 contain some background events for
    debugging detectors. Lots of missing ADCs and TDCs in
    TOFW. I was too tired to track this down but may well be
    HV. It seemed that the new DAQ code to read the 2nd ZDC
    ADC caused problems. At about 5am Konstantin set the
    DAQ back to the old code. Since then DAQ is OK and we
    get ZDC data for the 1st ADC. Blue ZDC seems to get
    activeated by beam. It was counting at 20Hz without
    beam. 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sat Jul 15 2000 - 08:02:51 EDT